Define an operator T in End(F^2) by T(x,y)= (y,0) Let
U = {(x,0) | x in F}. Show that
U is invariant under T and T |U is the 0 operator on U;
1
Expert's answer
2020-06-11T20:51:24-0400

Given that, "T" is an operator such that "T\in End(F^2)" and defined by

"T(x,y)= (y,0)"

And,

"U := \{(x,0) | x \in F\}"

Since, if "U" is invariant under "T" ,thus

"T(U)\subset U"

In this case, suppose "u\in U\implies u=(x,0)" for some "x\in F" ,Thus by definition

"T(u)=T(x,0)=(0,0)\in U \hspace{1cm}(\because 0\in U)\\
\implies T(U)\subset U"

Now, we have to show the restriction of "T" under "U" i.e

"T|_U:U\longrightarrow U"

is operator.

Clearly, for all

"v\in U\implies v\in F^2
\implies T|_U(v)=T(v)=0\\
\implies T|_U=0"

Hence we are done.